# On the Origin of Abundance Variations in the Milky Way’s High- Plateau

### v1

@@ 1 Introduction
Stars in the Milky Way (MW) disk follow two distinct sequences in the plane of [/Fe] vs. [Fe/H], with the “high-” and “low-” populations associated kinematically with the hot, thick disk and the cold, thin disk, respectively (Bensby et al., 2003). Thin disk stars are generally younger than thick disk stars of the same metallicity, and the lower (around solar) [/Fe] in the thin disk is understood as a consequence of greater Fe enrichment from time-delayed Type Ia supernovae (SNIa). The [/Fe] ratios of the high- population could reflect the yield ratios for pure core collapse supernova (CCSN) enrichment averaged over the stellar initial mass function (IMF), or they could represent a mix of SNIa and CCSN contributions but with a SNIa/CCSN ratio lower than that of the thin disk. In either case, it is more physically accurate to describe the “high-” and “low-” populations as “low-Ia” and “high-Ia,” respectively, which is the nomenclature we adopt in the rest of this paper. Using data from the SDSS APOGEE survey (Majewski et al., 2017), Bertran de Lis et al. (2016) estimate an intrinsic scatter of 0.04 dex in [O/Fe] at fixed [Fe/H] within the low-Ia population. Similarly, Vincenzo et al. (2021) estimate intrinsic scatter of 0.03-0.04 dex in the low-Ia population for [O/Fe], [Mg/Fe], [Si/Fe], and [Ca/Fe].
@@
In this paper, we investigate whether abundance scatter within the low-Ia population is driven primarily by variations in the SNIa/CCSN ratio or arises instead from other effects such as stochastic sampling of the supernova population. If there are significant variations of SNIa/CCSN within the low-Ia population, then elements that have a large SNIa contribution should vary together. For example, stars that are high in [Fe/Mg] should also be high in [Ni/Mg] and [Mn/Mg], where we have adopted Mg as our reference element because it is expected to arise entirely from CCSN. Conversely, elements whose production is dominated by CCSN should exhibit low scatter in [X/Mg] at fixed [Mg/H]. To implement this idea quantitatively, we adopt the two-process model (Weinberg et al., 2019, 2022; Griffith et al., 2019, 2022, 2024a; Sit et al., 2024), which describes a star’s multi-element abundances as the sum of enrichment from a prompt process associated with CCSN and a delayed process associated with SNIa. The two-process model has been successful at predicting the multi-element abundance pattern of stars in the MW disk to 0.020.05 dex accuracy, comparable to the observational uncertainties reported by large spectroscopic surveys. At each [Mg/H], the relative contribution of the two processes for a given element X is inferred from the gap in median [X/Mg] between the disk’s low-Ia and high-Ia populations. If abundance scatter in the low-Ia population comes mainly from SNIa/CCSN variation, then the two-process model should predict the element-by-element trends within this population, even though it is calibrated to the difference between low-Ia and high-Ia stars.
@@
We concentrate our analysis on low-Ia stars with , a range where the median trend of [Mg/Fe] with [Fe/H] is well described by a flat “plateau” at [Mg/Fe]. Many observational studies show this plateau continuing down to , with similar flat plateaus for [Si/Fe] and [Ca/Fe] (see, e.g., Kobayashi et al., 2020). It is tempting, therefore, to identify these plateau ratios with the ratios produced by CCSN alone. However, the star-to-star scatter and study-to-study variations become larger at . Furthermore, Conroy et al. (2022) find that in situ halo stars show a trend of declining [Mg/Fe] with increasing [Fe/H] for , rather than a flat plateau. They propose a model in which the true ratio from CCSN enrichment corresponds to [Mg/Fe], and the roughly flat trend of [Mg/Fe] in the thick disk reflects a balance between SNIa and CCSN enrichment during an epoch of accelerating star formation (see also Maoz & Graur, 2017; Chen et al., 2023).
@@
The two-process model recasts each star’s abundances into two amplitude parameters corresponding to the level of prompt (CCSN) and delayed (SNIa) contribution, and the element-by-element residuals between the observed abundances and the model predictions. Both the “process vectors” describing the relative element contributions and the inferred values of the star-by-star “process amplitudes” depend on the value of [Mg/Fe] that is assumed to represent pure CCSN enrichment (see Section 3 below). We originally hoped that our analysis could distinguish a scenario in which the thick disk plateau reflects CCSN only from a scenario in which SNIa enrichment is already substantial at . However, we show in Section 3.1 that the residual abundances from the two-process predictions are independent of the assumed [Mg/Fe]cc, even though the process amplitudes are not (see also Griffith et al. 2024a). Unfortunately, this means that our results do not shed light on the true level of [Mg/Fe]cc, though they do provide strong evidence for variation of the SNIa/CCSN ratio within the low-Ia population.
@@ 2 Data
Sit et al. (2024, hereafter S24) present corrected abundances for 288,789 stars from APOGEE DR17 (Abdurro’uf et al., 2022), where artificial trends in APOGEE abundances with stellar surface gravity were identified and removed. This calibration also removes trends with stellar effective temperature . In this work, we use a calibration sample and a smaller plateau sample, both selected from the full Sit et al. (2024) catalog. Our calibration sample is the same as that of S24, whose selection criteria are:
@@
The calibration sample consists of 151,564 stars from APOGEE DR17.
We also generate a smaller plateau sample of stars that occupy a well-populated region on the low-Ia plateau and is a subset of the calibration sample. We apply the following cuts in [Mg/Fe] and [Fe/H] to the calibration sample to define the plateau sample:
These cuts select the region with the highest density of plateau stars around the [Mg/Fe]cc = 0.3 value used in Weinberg et al. (2022) and S24. The plateau sample contains 10,391 stars. Figure 1 shows the calibration sample, divided into low-Ia and high-Ia populations, and the additional [Mg/Fe] and [Fe/H] cuts that define the plateau sample. The low-Ia population is defined as
and calculated using the raw ASPCAP values (i.e., prior to the calibration), the same as S24 and previous works (Weinberg et al., 2019, 2022).
We utilize 17 elements, also analyzed by S24, spanning a range of nucleosynthetic sources: O, Si, S, and Ca ( elements); C+N, Na, Al, and K (light odd-Z elements); Cr, Fe, Ni, V, Mn, Co, and Cu (iron-peak elements), and Ce (-process element). The combined element C+N accounts for the surface abundances of C and N changing as a star evolves up the red giant branch while the total number of nuclei remains nearly equal to the birth abundance. We suspect, but do not know, that most of the variation in [(C+N)/Mg] within our plateau sample comes from variation in birth [N/Mg], because CCSN are expected to be the primary sources of C while AGB stars make a large and metallicity-dependent contribution to N enrichment (see, e.g., Andrews et al., 2017; Rybizki et al., 2017; Johnson et al., 2023). S24 incorporated abundance measurements from the BAWLAS catalog (Hayes et al., 2022) into their analysis for O, C+N, S, Na, V, and Ce, but we use the ASPCAP measurements for this work to increase the sample size of available measurements, since BAWLAS used a higher SNR threshold of SNR 150. Cu is measured only by BAWLAS so we proceed with including it here while noting that only 2,156 out of 10,391 stars in our sample have a Cu measurement.
@@ 3 The Two-Process Model
The two-process model describes a star’s multi-element abundances as a sum of two contributions, associated with CCSN and SNIa, with amplitudes and . For each process, the relative contribution to different elements is taken to be universal, though it may depend on metallicity. These relative contributions reflect the population-averaged yields of CCSN and SNIa, where the population-average includes averages over the stellar IMF, binary properties, rotation, and other parameters that affect yields. In the notation of Weinberg et al. (2022, hereafter W22), the abundances predicted by the two-process model are
where the “process vectors” and are derived from the ensemble of stars and the amplitudes and are fit to each star individually. A star’s actual abundances may differ from the two-process prediction because additional processes contribute or because the CCSN and SNIa processes are not completely universal. The approach can be generalized to include additional processes (W22; Griffith et al. 2024a, hereafter G24), though these become more difficult to disentangle as the number of processes increases. More generally, one can think of the two processes as representing prompt and time-delayed nucleosynthesis contributions, and the two-process model is reasonably successful at predicting abundances of elements whose delayed contribution is expected to arise from asymptotic giant branch (AGB) stars rather than SNIa, such as C, N, Ce, Y, and Ba (Griffith et al. 2022; W22).
@@ 3.1 Process Vectors and Amplitudes
For present purposes, we would like to know how the predicted and residual abundances depend on the value [Fe/Mg] that is assumed to represent the iron-to-magnesium ratio from CCSN enrichment with no SNIa contribution. We follow the definition and equations of W22, and to simplify expressions we introduce the notation
@@
For stars in the plateau sample (see Section 2), we use S24 -calibrated abundances to derive and . Figure 3 shows how the distribution of values shifts to higher values as lower values are assumed. The overall distribution at fixed [Fe/Mg]cc is very similar between the Fe and Mg calculation (using Equations 7-8) and the six-element fit; the main notable difference is a slightly longer tail towards high for the six-element fit. The shift from [Fe/Mg] and [Fe/Mg] is smaller than [Fe/Mg] to [Fe/Mg], similarly to how the and sequences shift in Figure 2 for Ni, Mn, and Ce.
@@ 3.2 Residual Abundances and Linearity
We reduce the uncertainty in and from measurement uncertainties by using the six-element -minimization procedure, so the arguments given above are no longer exact. However, the star-by-star changes in , , and residual abundances between a two-element and six-element fit are small. By repeating our full procedure with different [Fe/Mg]cc choices we find that residual abundances are still nearly independent of the adopted value.
@@ 4.1 Variations Correlated with SNIa/CCSN Ratio
A plausible source of intrinsic abundance scatter in stars on the low-Ia plateau is remaining variation in the amount of SNIa to CCSN contribution. If this effect is dominant, we would expect that elements with significant SNIa contribution, such as the Fe-peak elements, should vary together, while elements with mostly CCSN contribution, like the elements, should exhibit little correlated variation. We separate the scatter from overall trends by defining the quantity [X/Mg]med, the difference between a star’s [X/Mg] and the median [X/Mg] of the plateau sample at a similar [Mg/H].
@@
Figure 7 more clearly demonstrates the definition of [X/Mg]med for example elements Fe, Si, and Mn. The median in [X/Mg] is calculated for the plateau sample (individual measurements shown in gray points, same as in Figure 6) in 0.1 dex wide bins; the median values are shown in the figure panels by the solid black line and effectively form a piecewise step function. Two example stars are highlighted in blue and red; observed ([X/Mg], [Mg/H]) points are indicated by the filled triangles, with the point facing upwards (downwards) if [X/Mg] is above (below) the median. For a given star, [X/Mg]med is the difference between its measured [X/Mg] and the median [X/Mg] calculated for its [Mg/H] bin, and [X/Mg]med is defined such that this quantity is positive if the star falls above the median. Thus, the length of the solid line connecting the black median to the triangle points is [X/Mg]med. One can also calculate [X/Mg]med for a predicted abundance, such as from the two-process model. These are shown for Si and Mn with the open triangles for the predicted [X/Mg] value and dotted lines for the predicted [X/Mg]med. For clarity, we define [X/Mg]med,obs as [X/Mg]med calculated from the observed data, and [X/Mg]med,2proc as [X/Mg]med calculated from two-process predicted abundances.
The two example stars in Figure 7 illustrate the predicted trend where variation is correlated in Fe-peak elements with significant SNIa contribution to their nucleosynthesis. The blue star falls below the median (i.e., has negative [X/Mg]med,obs) in both [Fe/Mg] and [Mn/Mg], and the red star falls above the median in both elements (i.e., has positive [X/Mg]med,obs). However, the variation in [Si/Mg] is minimal for the same stars, which is in line with Si being a CCSN element.
@@
We show this trend in [X/Mg]med between different elements for the full plateau sample in Figure 8. The slope of the [X/Mg]med vs. [Fe/Mg]med is noticeably steeper for Mn than for Si. The slope differences echo that seen in Figure 5. We could examine such correlations in the median differences between all the element pairs in our sample, but there is an important caveat in this test: the effect of measurement uncertainty, which adds scatter and can induce spurious correlations because Mg affects both x and y-axis quantities.
@@
In the two-process model fit to only Fe and Mg, the ratio of process amplitudes is a simple linear transformation of linear {Fe/Mg} (Equation 8, Figure 5). Our six-element fit gives a less noisy estimate of , so it is preferable to [Fe/Mg]med,obs when examining the trends with other elements’ [X/Mg]med,obs. An additional advantage of incorporating the two-process model is that we can use the predicted abundances to simulate how the elements should correlate if all intrinsic variation is explained by variation in SNIa/CCSN ratio and the only noise is Gaussian with the reported uncertainties.
@@
The slopes are similar between the observed data and two-process predictions for most elements, indicating that the two-process model predicts SNIa contribution on the plateau well. The two elements with the largest differences between their observed and predicted slopes are Ce and C+N. The slope calculated with [Ce/Mg]med,2proc is steeper than the slope calculated with [Ce/Mg]med,obs, indicating that the two-process model over-predicts the SNIa contribution to Ce. Because the delayed contribution to Ce probably reflects AGB enrichment rather than SNIa enrichment, it is not surprising that the linear trend inferred from the low-Ia and high-Ia medians does not apply within the low-Ia population (see Figure 5). Within our plateau sample, [Ce/Mg] exhibits little trend with , but large scatter. Conversely, the [(C+N)/Mg]med,2proc slope is shallower than the [(C+N)/Mg]med,obs slope. As with Ce, we expect the delayed contribution to C and N to come from AGB enrichment rather than SNIa enrichment, so it is not surprising that the linear trend between the low-Ia and high-Ia medians does not describe the behavior within the low-Ia population. However, it is notable that the prediction is too steep for Ce but too shallow for C+N. This different behavior could possibly arise from different metallicity dependence of C, N, and Ce yields, but we have no full explanation for these trends. The departures of C+N and Ce trends from those predicted based on the low-Ia and high-Ia median may provide clues to the physics of AGB enrichment in the early MW disk.
Other elements with small slope discrepancies include K, V, and Co, but these elements have large observational uncertainties, and non-Gaussian abundance errors or incorrect error estimates might affect the slopes. Finally, Ni also shows a small discrepancy between observed and predicted slopes. This is unexpected because Ni is a well-measured element used in the six-element two-process fit and is not known to have any nucleosynthetic contributions from processes outside of SNIa and CCSN. If the discrepancy is real, a possible explanation is that the Ni yields are higher at plateau metallicities.

### v2

@@ 1 Introduction
Stars in the Milky Way (MW) disk follow two distinct sequences in the plane of [/Fe] vs. [Fe/H], with the “high-” and “low-” populations associated (though not perfectly: see Hayden et al. 2017) with the kinematic hot, thick disk and the cold, thin disk, respectively. Low- stars are generally younger than high- stars of the same metallicity, and the lower (around solar) [/Fe] is understood as a consequence of greater Fe enrichment from time-delayed Type Ia supernovae (SNIa) (e.g., Matteucci & Greggio, 1986; McWilliam, 1997; Fuhrmann, 1998; Bensby et al., 2003; Adibekyan et al., 2012). While the “knee” where the [/Fe] ratio begins to decrease with increasing metallicity is a signature of SNIa enrichment in the thick disk (e.g., Feltzing et al., 2003), a common assumption is that the [/Fe] ratios in the high- populationat metallicities below the knee, where the [/Fe] trend is mostly flat, reflect the yield ratios for pure core collapse supernova (CCSN) enrichment, averaged over the stellar initial mass function (IMF) and other progenitor properties like rotation and binarity. Alternatively, the higher [/Fe] ratio could represent a mix of SNIa and CCSN contributions, but with a SNIa/CCSN ratio lower than that of the low- population. In either case, it is more physically accurate to describe the “high-” and “low-” populations as “low-Ia” and “high-Ia,” respectively, which is the nomenclature we adopt in the rest of this paper. Using data from the SDSS APOGEE survey (Majewski et al., 2017), Bertran de Lis et al. (2016) estimate an intrinsic scatter of 0.04 dex in [O/Fe] at fixed [Fe/H] within the low-Ia population. Similarly, Vincenzo et al. (2021) estimate intrinsic scatter of 0.03-0.04 dex in the low-Ia population for [O/Fe], [Mg/Fe], [Si/Fe], and [Ca/Fe].
@@
In this paper, we investigate the abundance scatter of 17 elements within the low-Ia population. One key question is whether abundance scatter within the low-Ia population is driven primarily by variations in the SNIa/CCSN ratio or arises instead from other effects such as stochastic sampling of the supernova population. If there are significant variations of SNIa/CCSN within the low-Ia population, then elements that have a large SNIa contribution should vary together. For example, stars that are high in [Fe/Mg] should also be high in [Ni/Mg] and [Mn/Mg], where we have adopted Mg as our reference element because it is expected to arise entirely from CCSN with little metallicity dependence (see, e.g., Nomoto et al., 2013; Andrews et al., 2017, and references therein). Conversely, elements whose production is dominated by CCSN should exhibit low scatter in [X/Mg] at fixed [Mg/H].
To implement this idea quantitatively, we adopt the two-process model (Weinberg et al., 2019, 2022; Griffith et al., 2019, 2022, 2024a; Sit et al., 2024), which describes a star’s multi-element abundances as the sum of enrichment from a prompt process associated with CCSN and a delayed process associated with SNIa. The two-process model has been successful at predicting the multi-element abundance pattern of stars in the MW disk to 0.020.05 dex accuracy, comparable to the observational uncertainties reported by large spectroscopic surveys. At each [Mg/H], the relative contribution of the two processes for a given element X is inferred from the gap in median [X/Mg] between the disk’s low-Ia and high-Ia populations. If abundance scatter in the low-Ia population comes mainly from SNIa/CCSN variation, then the two-process model should predict the element-by-element trends within this population, even though it is calibrated to the difference between low-Ia and high-Ia stars. However, the two-process model is limited by its namesake two processes, SNIa and CCSN, so deviations can be indicative of other nucleosynthetic pathways, such as enrichment from asymptotic giant branch (AGB) stars (Griffith et al., 2022; Weinberg et al., 2022). Therefore, we can also investigate the contribution of any additional processes to the abundance scatter in the low-Ia plateau.
@@
We concentrate our analysis on low-Ia stars with , a range where the median trend of [Mg/Fe] with [Mg/H] is well described by a flat “plateau” at [Mg/Fe]. Many observational studies show this plateau continuing down to , with similar flat plateaus for [Si/Fe] and [Ca/Fe] (see, e.g., Kobayashi et al., 2020a). It is tempting, therefore, to identify these plateau ratios with the ratios produced by CCSN alone. However, the star-to-star scatter and study-to-study variations become larger at . Furthermore, Conroy et al. (2022) find that in situ halo stars show a trend of declining [Mg/Fe] with increasing [Fe/H] for , rather than a flat plateau. They propose a model in which the true ratio from CCSN enrichment corresponds to [Mg/Fe], and the roughly flat trend of [Mg/Fe] in the thick disk reflects a balance between SNIa and CCSN enrichment during an epoch of accelerating star formation (see also Maoz & Graur, 2017; Chen et al., 2023).
@@
The two-process model recasts each star’s abundances into two amplitude parameters corresponding to the level of prompt (CCSN) and delayed (SNIa) contribution, and the element-by-element residuals between the observed abundances and the model predictions. Both the “process vectors” describing the relative element contributions and the inferred values of the star-by-star “process amplitudes” depend on the value of [Mg/Fe] that is assumed to represent pure CCSN enrichment (see Section 3 below). We originally hoped that our analysis could distinguish a scenario in which the low-Ia plateau reflects CCSN only from a scenario in which SNIa enrichment is already substantial at . However, we show in Section 3.1 that the residual abundances from the two-process predictions are independent of the assumed [Mg/Fe]cc, even though the process amplitudes are not (see also Griffith et al. 2024a). Unfortunately, this means that our results do not shed light on the true level of [Mg/Fe]cc, though they do provide strong evidence for variation of the SNIa/CCSN ratio within the low-Ia population.
@@ 2 Data
Systematic trends in abundance measurements with and can cause changes in median abundance trends (Griffith et al., 2021a) or, more importantly for this study, add correlated scatter between elements (Weinberg et al., 2022). These trends are most likely an artifact of the spectral fitting process, as a star’s changes during its post-main sequence lifetime, but the abundances of most elements should stay the same. Sit et al. (2024, hereafter S24) present corrected abundances for 288,789 stars from APOGEE DR17 (Abdurro’uf et al., 2022), where artificial trends in APOGEE abundances with stellar surface gravity were identified and removed. This calibration also removes trends with stellar effective temperature .
In this work, we use a calibration sample and a smaller plateau sample, both selected from the full S24 catalog. Our calibration sample is the same as that of S24, whose selection criteria are:
@@
The calibration sample consists of 149,904 stars from APOGEE DR17.
We also generate a smaller plateau sample of stars that occupy a well-populated region on the low-Ia plateau and is a subset of the calibration sample. We apply the following cuts in [Fe/Mg] and [Mg/H] to the calibration sample to define the plateau sample:
These cuts select the region with the highest density of plateau stars around the [Mg/Fe]cc = 0.3 value used in Weinberg et al. (2022) and S24. The plateau sample contains 10,262 stars.
Figure 1 shows the calibration sample, divided into low-Ia and high-Ia populations, and the additional [Mg/Fe] and [Fe/H] cuts that define the plateau sample. The low-Ia population is defined as
For this division between the low-Ia and high-Ia populations only, we use the raw ASPCAP values (i.e., prior to the calibration). This maintains consistency with S24 in the median [X/Mg] sequences of each population from which the “process vectors” of the two-process model are derived (see Section 3), so that we better isolate the effect of varying the assumed core-collapse [Fe/Mg] ratio in the two-process model in Section 3. After accounting for flat offsets applied during calibration (see S24 for details), applying the calibrations would change the high/low-Ia classification of only 1% of the calibration sample, and only for stars near the division boundary (none of which are in the plateau sample).
We utilize 17 elements, also analyzed by S24, spanning a range of nucleosynthetic sources: O, Si, S, and Ca ( elements); C+N, Na, Al, and K (light odd-Z elements); Cr, Fe, Ni, V, Mn, Co, and Cu (iron-peak elements), and Ce (-process element). The combined element C+N accounts for the surface abundances of C and N changing as a star evolves up the red giant branch while the total number of nuclei remains nearly equal to the birth abundance. S24 incorporated abundance measurements from the BAWLAS catalog (Hayes et al., 2022) into their analysis for O, C+N, S, Na, V, and Ce, but we use the ASPCAP measurements for this work to increase the sample size of available measurements, since BAWLAS used a higher SNR threshold of SNR 150. Cu is on the sharply falling edge of the iron peak (“Fe-cliff”) and though its astrophysical origins are debated, literature suggests more significant contributions from massive stars (e.g., Woosley & Weaver, 1995; Romano & Matteucci, 2007; Kobayashi et al., 2020a), so it may have different nucleosynthesis patterns from the other Fe-peak elements. Therefore, despite being measured only by BAWLAS, we proceed with including Cu in this study, but note that only 2,277 out of 10,262 stars in our sample have a Cu measurement.
As part of our analysis in this paper, we use the reported measurement errors from ASPCAP (or BAWLAS, for Cu). The measurement uncertainties in DR17 are based on the process for DR16, where a function of , [M/H], and SNR is fit to the deviations in measured abundances from repeat observations (see Section 5.4 of Jönsson et al. 2020 for more details). A similar procedure using repeat measurements is used for the empirical BAWLAS uncertainties (see Section 4.7.2 in Hayes et al. 2022) we use for Cu.
@@ 3 The Two-Process Model
The two-process model is a semi-empirical nucleosynthesis model that describes a star’s multi-element abundances as a sum of two contributions, associated with CCSN and SNIa, with amplitudes and . The processes are anchored in observed median [X/Mg] trends with metallicity, which have been observed to be nearly independent of location across the Galactic disk provided the low-Ia and high-Ia populations are separated (Weinberg et al., 2019). By interpreting these spatially invariant trends as driven by nucleosynthetic yields rather than spatially-varying quantities such as the star formation history, the relative contribution to different elements is taken to be universal for each process (though it may depend on metallicity). These relative contributions reflect the population-averaged yields of CCSN and SNIa, where the population-average includes averages over the stellar IMF, binary properties, rotation, and other parameters that affect yields. In the notation of Weinberg et al. (2022, hereafter W22), the abundances predicted by the two-process model are
where the “process vectors” and are derived from the ensemble of stars and the amplitudes and are fit to each star individually.
The two-process model can successfully predict stellar abundances in the MW disk to 0.02-0.05 dex on average (e.g., Weinberg et al. 2019; Griffith et al. 2019; W22), and similar accuracy has been achieved with other two-parameter models conditioned on age and [Fe/H] (Ness et al., 2019) or and Fe abundances (Ness et al., 2022; Ratcliffe & Ness, 2023). However, a star’s actual abundances may differ from the two-process prediction because additional processes contribute or because the CCSN and SNIa processes are not completely universal. The approach can be generalized to include additional processes (W22; Griffith et al. 2024a, hereafter G24), though these become more difficult to disentangle as the number of processes increases. More generally, one can think of the two processes as representing prompt and time-delayed nucleosynthesis contributions, and the two-process model is reasonably successful at predicting abundances of elements whose delayed contribution is expected to arise from asymptotic giant branch (AGB) stars rather than SNIa, such as C, N, Ce, Y, and Ba (Griffith et al. 2022; W22).
@@ 3.1 Process Vectors and Amplitudes
For present purposes, we would like to know how the predicted and residual abundances depend on the value [Fe/Mg]cc that is assumed to represent the iron-to-magnesium ratio from CCSN enrichment with no SNIa contribution. We test assumptions of [Fe/Mg]cc values spaced at 0.1 dex intervals between [Fe/Mg], the fiducial assumption from W22 and S24, and [Fe/Mg], approximately the value estimated by Conroy et al. (2022). We follow the definition and equations of W22, and to simplify expressions we introduce the notation
@@
For stars in the plateau sample (see Section 2), we use S24 -calibrated abundances to derive and . Figure 3 shows how the distribution of values shifts to higher values as lower values are assumed. The overall distribution at fixed [Fe/Mg]cc is very similar between the Fe and Mg calculation (using Equations 7-8) and the six-element fit; the main notable difference is a slightly longer tail towards high for the six-element fit. The shift from [Fe/Mg] and [Fe/Mg] is smaller than [Fe/Mg] to [Fe/Mg], similarly to how the and sequences shift in Figure 2 for Ni, Mn, and Ce. Additionally, there are no variations in correlated with guiding radius or maximum height from the midplane within the plateau sample. These values are taken from the astroNN catalog for DR17222https://www.sdss4.org/dr17/data_access/value-added-catalogs/?vac_id=the-astronn-catalog-of-abundances,-distances,-and-ages-for-apogee-dr17-stars, which uses Gaia eDR3 for distances via Leung & Bovy (2019) and derives orbital parameters via Mackereth & Bovy (2018).
@@ 3.2 Residual Abundances and Linearity
We reduce the uncertainty in and from measurement uncertainties by using the six-element -minimization procedure, so the arguments given above are no longer exact. However, the star-by-star changes in , , and residual abundances between a two-element and six-element fit are small. By repeating our full procedure with different [Fe/Mg]cc choices we find that residual abundances are still nearly independent of the adopted value. Therefore, the residual abundances (though not the two-process parameters) presented in the S24 catalog remain valid over all assumptions of [Fe/Mg]cc.
@@ 4.1 Variations Correlated with SNIa/CCSN Ratio
A plausible source of intrinsic abundance scatter in stars on the low-Ia plateau is remaining variation in the amount of SNIa to CCSN contribution. If this effect is dominant, we would expect that elements with significant SNIa contribution, such as the Fe-peak elements, should vary together, while elements with mostly CCSN contribution, like the elements, should exhibit little correlated variation.
@@
We separate the scatter from overall trends by defining the quantity [X/Mg]med, the difference between a star’s [X/Mg] and the median [X/Mg] of the plateau sample at a similar [Mg/H].The median in [X/Mg] is calculated for the plateau sample in 0.1 dex wide bins centered at [Mg/H] = , and . The medians are shown by the solid black line in each panel in Figure 6. For a given star, [X/Mg]med is the difference between its measured [X/Mg] and the median [X/Mg] calculated for its [Mg/H] bin, thus removing overall trends with metallicity. [X/Mg]med is defined such that this quantity is positive if the star falls above the median. One can also calculate [X/Mg]med for a predicted abundance, such as from the two-process model. For clarity, we define [X/Mg]med,obs as [X/Mg]med calculated from the observed data, and [X/Mg]med,2proc as [X/Mg]med calculated from two-process predicted abundances.
@@
We show an example of the expected trend in [X/Mg]med between different elements for the full plateau sample in Figure 7. The slope of the [X/Mg]med vs. [Fe/Mg]med is noticeably steeper for Mn than for Si, consistent with with Mn having significantly more SNIa contribution than Si. The slope differences echo that seen in Figure 5. We could examine such correlations in the median differences between all the element pairs in our sample, but there is an important caveat in this test: the effect of measurement uncertainty, which adds scatter and can induce spurious correlations because Mg affects both x and y-axis quantities.
@@
In the two-process model fit to only Fe and Mg, the ratio of process amplitudes is a simple linear transformation of linear {Fe/Mg} (Equation 8, Figure 5). Our six-element fit gives a less noisy estimate of , so it is preferable to [Fe/Mg]med,obs when examining the trends with other elements’ [X/Mg]med,obs.In a small fraction of stars, the six-element fit results in a best-fit that is outside the bounds of the calculated from the linear transformation of Fe/Mg (which is set by our sample selection), resulting in a more extended distribution compared to [Fe/Mg]med,obs. An additional advantage of incorporating the two-process model is that we can use the predicted abundances to simulate how the elements should correlate if all intrinsic variation is explained by variation in SNIa/CCSN ratio and the only noise is Gaussian with the reported uncertainties.
@@
The slopes are similar between the observed data and two-process predictions for most elements, indicating that the two-process model predicts SNIa contribution on the plateau well. The two elements with the largest differences between their observed and predicted slopes are Ce and C+N. The slope calculated with [Ce/Mg]med,2proc is steeper than the slope calculated with [Ce/Mg]med,obs, indicating that the two-process model over-predicts the SNIa contribution to Ce. Because the delayed contribution to Ce probably reflects AGB enrichment rather than SNIa enrichment, it is not surprising that the linear trend inferred from the low-Ia and high-Ia medians does not apply within the low-Ia population (see Figure 5). Within our plateau sample, [Ce/Mg] exhibits little trend with , but large scatter.
Conversely, the [(C+N)/Mg]med,2proc slope is shallower than the [(C+N)/Mg]med,obs slope. As with Ce, we expect the delayed contribution to C and N to come from AGB enrichment rather than SNIa enrichment, so it is not surprising that the linear trend between the low-Ia and high-Ia medians does not describe the behavior within the low-Ia population. However, it is notable that the prediction is too steep for Ce but too shallow for C+N. This different behavior could possibly arise from different metallicity dependence of C, N, and Ce yields, but we have no full explanation for these trends. We suspect, but do not know, that most of the variation in [(C+N)/Mg] within our plateau sample comes from variation in birth [N/Mg], because CCSN are expected to be the primary sources of C while AGB stars make a large and metallicity-dependent contribution to N enrichment (see, e.g., Andrews et al., 2017; Rybizki et al., 2017; Johnson et al., 2023). The departures of C+N and Ce trends from those predicted based on the low-Ia and high-Ia median may provide clues to the physics of AGB enrichment in the early MW disk.
Other elements with small slope discrepancies include K, V, and Co, but these elements have large observational uncertainties, and non-Gaussian abundance errors or incorrect error estimates might affect the slopes. Finally, Ni also shows a small discrepancy between observed and predicted slopes. This is unexpected because Ni is considered a well-measured element in ASPCAP that we used in the six-element two-process fit and is not known to have any nucleosynthetic contributions from processes outside of SNIa and CCSN. When NLTE effects are accounted for, the ratio of Ni to Fe increases as metallicity decreases at subsolar metallicities (Eitner et al., 2023); this could possibly be a related signature. If the discrepancy is real, a possible explanation of the steeper Ni slope is that the population-averaged Ni yields are higher at plateau metallicities. A potential cause of different yields may be differences in the SNIa progenitor landscape. Ni, Mn, and Cr yields are known to be different between sub-Chandrasekhar and near-Chandrasekhar mass SNIa (Kobayashi et al., 2020b). Therefore, if the ratio of sub-Chandrasekhar to near-Chandrasekhar white dwarfs is different at the time of plateau formation compared to the current MW disk, this would result in different SNIa yields on the plateau even at same metallicity. However, as the Mn (and to a lesser extent, Cr) yields are also affected in this scenario, we would expect to also see a discrepant slope in Mn, but the difference between the [Mn/Mg]med,obs and [Mn/Mg]med,2proc slopes is smaller than for Ni.